You decide to go for a walk on the tundra to enjoy the Northern Lights. If you walk 2.0 km [W] and then 3.2 km [N 20 W] and then 1.5 km [S], what is your total displacement?




1
Expert's answer
2021-02-22T10:21:50-0500

Let's find "x" and "y" components of the resulting displacement:


"d_x=2.0\ km\cdot cos180^{\circ}+3.2\ km\cdot cos160^{\circ}=-5.0\ km,""d_y=3.2\ km\cdot sin160^{\circ}+1.5\ km\cdot sin270^{\circ}=-0.405\ km."

Then, the total displacement can be found from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(-5.0\ km)^2+(-0.405\ km)^2}=5.02\ km."

We can find the angle as follows:


"\theta=sin^{-1}(\dfrac{d_y}{d})=sin^{-1}(\dfrac{-0.405\ km}{5.02\ km})=-4.63^{\circ}."


The sign minus means that resultant displacement has direction "4.63^{\circ} S\ of\ W".

Therefore, the total displacement "d" has magnitude 5.02 km and direction of "4.63^{\circ} S\ of\ W".